OpenStudy (anonymous):

If x is an odd integer, y and z are whole numbers, and x(y+z)+x is odd, which of the following must be true?

OpenStudy (anonymous):

z is even z is odd y and z are either both odd or both even if y is odd, then z is even if y is even, then z is odd

OpenStudy (anonymous):

Well we know the product of an odd and an even gives an even, and the product of an odd with an odd gives us an odd. We know that the sum of two odds is even, and the sum of an odd and an even gives us an odd.

OpenStudy (anonymous):

Since \(x\) is odd and \(x+x(y+z)\) is also odd we can infer \(x(y+z)\) must then be even.

OpenStudy (anonymous):

y and z are either both odd or both even.

OpenStudy (anonymous):

that was a joke

OpenStudy (anonymous):

If \(x(y+z)\) is even, since we know \(x\) is odd, we can infer \(y+z\) must be even. When is the sum of two numbers even? Well when either both are odd or both are even! Therefore we know \(y,z\) are either both odd or both even.

OpenStudy (anonymous):

i mean that wasn't a joke

OpenStudy (loser66):

@oldrin.bataku I have question: x(y+z) +x = x((y+z+1) x odd * whatever = odd, right? so the answer is not one of the options

OpenStudy (anonymous):

no; an odd multiplied with an even is even

OpenStudy (loser66):

yes, my bad.

OpenStudy (loser66):

so, y +z +1 must be odd y +z must be even iff y and z both odd or both even, right?

OpenStudy (anonymous):

Consider an odd number \(2m+1\) for integer \(m\) and an even number \(2n\) for integer \(n\):$$2n(2m+1)=2n\cdot2m+2n=4nm+2n=2(2nm+n)$$ which is divisible by \(2\) and is therefore even.

OpenStudy (loser66):

another question: we discuss and give out the answer only, does it a real help?

OpenStudy (anonymous):

@Loser66 if \(y+z+1\) is odd then \(y+z\) must be even. If \(y+z\) is even then \(y,z\) must either be even or odd.

OpenStudy (anonymous):

Consider arithmetic \(\mod 2\); an even number is congruent to \(0\) and an odd number congruent to \(1\). We have \(0+0=0\pmod2\) and \(1+1=0\pmod2\)

OpenStudy (loser66):

like it. it's more professional,

OpenStudy (anonymous):

:-p and well I mean our discussion is public! hopefully OP will read our replies